# BFS-Prover: Scalable Best-First Tree Search for LLM-based Automatic Theorem Proving

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have spurred growing interest in automatic theorem proving using Lean4, where effective tree search methods are crucial for navigating the underlying large proof search spaces. While the existing approaches primarily rely on value functions and/or Monte Carlo Tree Search (MCTS), the potential of simpler methods like Best-First Tree Search (BFS) remains underexplored. In this paper, we investigate whether BFS can achieve competitive performance in large-scale theorem proving tasks. We present BFS-Prover, a scalable expert iteration framework, featuring three key innovations. First, we implement strategic data filtering at each expert iteration round, excluding problems solvable via beam search node expansion to focus on harder cases. Second, we improve the sample efficiency of BFS through Direct Preference Optimization (DPO) applied to state-tactic pairs automatically annotated with compiler error feedback, refining the LLM’s policy to prioritize productive expansions. Third, we employ length normalization in BFS to encourage exploration of deeper proof paths. BFS-Prover achieves a state-of-the-art score of $72.95\%$ on the MiniF2F test set and therefore challenges the perceived necessity of complex tree search methods, demonstrating that BFS can achieve competitive performance when properly scaled. To facilitate further research and development in this area, we have open-sourced our model at https://huggingface.co/bytedance-research/BFS-Prover.
@@ 1 Introduction
Automatic theorem proving (ATP) in formal languages has recently become a critical benchmark for evaluating the reasoning capabilities of large language models (LLMs). By encoding mathematical problems into formal systems like Lean4, ATP enables the generation of machine-verified proofs for complex mathematical propositions, ensuring logical correctness [12, 16]. Despite the remarkable success of LLMs in natural language mathematics and reasoning tasks [9, 13], theorem proving in formal languages presents unique challenges [28, 23, 6, 22]. Unlike informal reasoning, formal systems require strict adherence to syntax and semantics, as well as the ability to generate valid steps within a highly constrained formal framework. Additionally, the action (tactic) space in ATP is vast, as each proof state can lead to numerous potential tactics, making the search process for valid proofs computationally intensive [15].
Tree search algorithms are fundamental to ATP, allowing policy models to navigate large and complex proof space efficiently [16]. Among these methods, Monte Carlo Tree Search (MCTS) [4] has gained popularity due to its ability to balance exploration and exploitation using value functions (critic models) or intrinsic rewards [3, 18]. MCTS has demonstrated remarkable success in AlphaZero-style frameworks for games like chess and Go [19], where the underlying state spaces are vast but terminal states are well-defined. However, applying MCTS and/or value functions to ATP comes with distinct complications. While games have clear winning and losing conditions, proof search lacks such definitive end states: a proof attempt can theoretically continue indefinitely until a proof or counterexample is found, making it challenging to assess intermediate progress [5, 8]. Additionally, ATP involves a much larger and more dynamic branching factor, sparse and delayed feedback, and an open-ended reasoning process. These differences highlight the unique demands of ATP and advocate the need for specialized adaptations of search methods to address its complexities.
Best-First Tree Search (BFS) [14] offers a simpler and more lightweight alternative to MCTS by prioritizing expansions based on accumulated log probabilities from the current node to the root. While its straightforward nature and computational simplicity make it attractive, BFS is often considered suboptimal for theorem proving in the existing literature [24, 11, 26] due to the following hypothesis:
@@ 1.1 Our Contributions
  • Expert iteration with self-filtering. We develop an expert iteration [1] framework that strategically filters out problems solvable by beam search [20] node expansions in each round. This filtering is critical as it directs training data accumulation towards harder theorems. As expert iteration progresses, the policy LLM continuously improves, learning a diverse range of tactics and deeper proofs via BFS.
  • Direct preference optimization (DPO) from compiler feedback. We use DPO [17] to refine the policy LLM by leveraging preference pairs naturally generated during tree search. From a given proof state, each preference pair consists of a positive tactic, which lies on the proof path, and a negative tactic, which leads to a Lean compiler error. DPO sharpens policy’s distribution, enabling it to avoid unproductive tactics and thereby improving the sample efficiency of BFS.
@@
  • Empirical results on MiniF2F. BFS-Prover achieves an accumulative score of 72.95% on the MiniF2F test set, surpassing all state-of-the-art theorem provers in the literature, including DeepSeek-Prover-V1.5 [26], InternLM2.5-StepProver [24], and HunyuanProver [11]. This result underscores that BFS-Prover can scale to competitive levels in ATP while maintaining a lightweight design that avoids the complexities of MCTS and value functions.
Roadmap. The rest of this paper is organized as follows. Section 2 gives an overview of the BFS-Prover system, detailing the expert iteration framework, data filtering, DPO for policy refinement, and length normalization in BFS. Section 3 describes practical implementation and experimental results on the MiniF2F benchmark against leading prover systems. Section 4 concludes the paper.
@@ 2.1 Lean4 Environment and Policy LLM
We adopt LeanDojo [29] as an interactive Python interface for Lean4 integration within BFS-Prover. It transforms Lean4 into a gym-like environment [2], facilitating interaction between the policy LLM and the formal proof assistant. Specifically, LeanDojo manages state transitions by executing tactics generated by the policy LLM in the Lean4 compiler. If a tactic cannot be executed, LeanDojo captures and returns the corresponding error messages, providing critical feedback for DPO to refine the policy LLM.
@@ 2.3 Expert Iteration
    Direct Preference Optimization (DPO): Instead of re-training the policy LLM using SFT, DPO can be applied as an alternative method to refine the current policy model by leveraging the on-policy Lean error data collected in step 2. Along the proof path, certain generated tactics, expanded from a proof state but not part of the verified proof path, result in Lean compiler errors. These erroneous tactics naturally serve as negative examples when compared to their corresponding valid proof tactics, forming preference tuples. Specifically, for a state $s$ on a proof path, we may form pairs $(a_{w},a_{l})$ where $a_{w}$ is the tactic on the proof path and $a_{l}$ is an error-causing tactic expanded from $s$ if exists; see Fig. 1 for an illustration. The DPO loss $\mathcal{L}_{\text{DPO}}(\theta)$ [17] is then defined as
    $-\mathbb{E}_{(s,a_{w},a_{l})}\!\left[\log\sigma(\beta(r_{\theta}(s,a_{w})-r_{\theta}(s,a_{l})))\right]\!,$
@@ 3 Practical Implementation and Benchmark Results
In this section, we discuss practical implementation details of the BFS-Prover system and present its benchmark results on MiniF2F test.
@@ 3.1 Model, Data, and Training Setup
Base model and initial training data. For illustration purposes, we use Qwen2.5-Math-7B [27] as the base model for fine-tuning the policy LLM in BFS-Prover. In order to properly initialize the expert iteration process, we leverage proof data extracted by LeanDojo [29] from Mathlib [12], which serves as the cold-start dataset. As expert iteration progresses, we further incorporate state-tactic datasets from Lean-Github [25], a compilation of Lean4 repositories from GitHub, and Lean-Workbook [30], which focuses on Olympia-level algebra and analysis. These datasets span a diverse range of formal mathematics topics, equipping the policy LLM with the foundational skills necessary for effective tactic generation and proof navigation.
Formal statement corpus. To construct our expert iteration data corpus, we autoformalize the NuminaMath-CoT dataset [10] using proprietary tools. We augment this with unproven theorems from Mathlib and formal statements from Lean-Workbook. The resulting corpus comprises approximately 900,000 formal mathematical statements without proofs, providing a comprehensive foundation for expert iteration.
@@
BFS configuration in expert iteration. We set the length normalization parameter $\alpha$ to $0.0$ throughout, to minimize inductive bias in expert iteration. During the beam search filtering stage, we utilize a beam width of $32$ to identify easily solvable theorems. In the subsequent data collection phase, we employ temperature-based sampling with a temperature of $1.0$, nucleus sampling parameter $1.0$, and a sampling width of $2$, $4$, or $8$ to explore diverse proof paths.
@@ 3.2 Distributed Best-First Search Infrastructure
  • GPU-Based Policy LLM Pool: Each local machine deploys 8 instances of our 7B policy LLM, each powered by an asynchronous vLLM engine [7] running on a dedicated A100 GPU. These instances form a shared pool that processes concurrent tactic generation requests.
@@
Figure 3: Sample evolution of tactic token length distribution across expert iteration rounds. The plot shows the log-scaled percentage of tactics falling into different token length bins. We have two observations: (1) robustness against mode collapse during expert iteration, evidenced by the maintained diversity across token lengths, and (2) a gradual but meaningful shift from an emphasis on very short tactics (111110101010 tokens, decreasing by about $2.3\%$) to a higher prevalence of moderate-length tactics (1111111150505050 tokens, with incremental increases ranging from approximately $1.27\%$ to $0.06\%$). This evolution reflects the improved ability of BFS-Prover to generate more complex and informative tactics while still retaining the capacity to produce concise ones when needed.
@@ 3.3 Distribution Shift in Expert Iteration
Tactic level. Beyond proof-level evolution, we also observe interesting distributional shifts at the tactic level throughout expert iteration; see Fig. 3. Notably, the policy LLM in the BFS-Prover system maintains a diverse range of tactic lengths without collapsing to a narrow distribution - a common failure mode in reinforcement learning where models tend to converge to a limited set of high-reward actions [21]. Instead, we observe a moderate but meaningful shift from very simple tactics (1-10 tokens) towards more commonly used tactical patterns (11-50 tokens). This shift suggests that through expert iteration, the policy LLM in BFS-Prover learns to generate more sophisticated tactics while maintaining the capability to use simpler ones when appropriate. The preservation of tactic diversity is crucial for effective theorem proving, as different proof states require different levels of tactical complexity, from simple term rewriting to complex algebraic manipulations.
@@ 3.4 Results on MiniF2F
In this subsection, we discuss BFS-Prover’s performance on the MiniF2F test benchmark [31] that is a widely recognized dataset for evaluating formal mathematics systems. MiniF2F is consisted of a diverse collection of formalized competition-level mathematics problems. The policy LLM checkpoint for tactic generation used for evaluation was obtained by performing SFT on all state-tactic pairs accumulated through $10$ rounds of expert iteration pipeline of BFS-Prover, followed by an additional round of DPO refinement using Lean error signals as described in Section 2.3.
@@ 3.4.1 Comparison with the State of the Art
We now compare BFS-Prover developed in this work against leading theorem provers in the literature, including DeepSeek-Prover-V1.5 [26], InternLM2.5-StepProver [24], and HunyuanProver [11].
@@
Accumulative results. In order to maximize potential proof discovery, we conduct BFS searches under multiple hyperparameter configurations and aggregate the successful proofs. Specifically, we vary the BFS length normalization factor $\alpha$ defined in (1) over $\{0.0,0.5,1.0\}$ to execute different exploration and exploitation trade-offs, while other parameter settings remain unchanged. This diverse search strategy allows BFS-Prover to accumulatively achieve a score of 72.95% on the MiniF2F test, setting a new state-of-the-art result in the formal theorem proving literature.
@@
DeepSeek-Prover-V1.5 [26] No MCTS $32\times 16\times 400$ 63.5%
@@
InternLM2.5-StepProver [24] Yes BFS $256\times 32\times 600$ 65.9%
HunyuanProver [11] Yes BFS $600\times 8\times 400$ 68.4%
@@ 4 Conclusion and Discussion
While BFS-Prover demonstrates strong performance in ATP, several limitations should be acknowledged, particularly regarding model size considerations. Our current implementation relies on a relatively small 7B parameter policy LLM, which may limit the system’s ability to learn and utilize more sophisticated mathematical reasoning patterns. While larger models (e.g., 32B or 70B parameters) could potentially capture more complex mathematical insights and generate more nuanced tactics, they would introduce significant computational challenges for both training and inference in the context of tree search. This trade-off becomes particularly evident when considering that larger models typically require more GPU memory and have longer inference latency, which could significantly reduce the number of states that can be explored within a fixed time budget. Additionally, complex mathematical proofs can generate extensive state descriptions that may exceed the practical context window of a 7B model, potentially causing the model to miss crucial information needed for generating appropriate tactics.
Author Contributions
Ran Xin conceptualized and led the project, designed and implemented the core components of the prover system developed in this work.
Chenguang Xi developed the autoformalization data pipeline for expert iteration.
Jie Yang provided critical support for the in-house distributed infrastructure.
Feng Chen contributed to the initial development of the prover system and explored various algorithmic approaches including value functions, MCTS, and DPO.
Hang Wu contributed to project conceptualization and provided valuable insights into the application of tree search and DPO in automatic theorem proving.
Xia Xiao, Yifan Sun, and Shen Zheng contributed to the development of in-house infrastructure and optimization of training parameters.
Kai Shen supervised the project and provided strategic direction.
Acknowledgements
We are grateful to Jing Su, Kaibo Liu, and Yuyu Zhang from the Seed Foundation Code team for their support throughout this project. We thank Jianqiao Lu, Chen Zheng, and Xun Zhou from the Seed Foundation Model team for useful discussion in the early stage of this project.
Appendix: Lean4 Proofs for IMO Problems Found by BFS-Prover
We present Lean4 proofs found by BFS-Prover for several IMO problems in MinF2F test. These examples demonstrate that BFS-Prover can handle complex mathematical reasoning in formal systems, including number theory, inequalities, and geometric relationships. See Fig. 5, 6, 7, 8, and 9 for detailed proofs.
@@ Acknowledgements
Figure 6: Lean4 proof of IMO-1964-P2 found by BFS-Prover.
@@
     linarith [Real.sqrt_nonneg (1 + 2 * x)]
     ring_nf at h2 |-
     field_simp at h2
     contrapose! h2
     rw [le_div_iff]
     swap

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have spurred growing interest in automatic theorem proving using Lean4, where effective tree search methods are crucial for navigating the underlying large proof search spaces. While the existing approaches primarily rely on value functions and/or Monte Carlo Tree Search (MCTS), the potential of simpler methods like Best-First Tree Search (BFS) remains underexplored. In this paper, we investigate whether BFS can achieve competitive performance in large-scale theorem proving tasks. We present BFS-Prover, a scalable expert iteration framework, featuring three key innovations. First, we implement strategic data filtering at each expert iteration round, excluding problems solvable via beam search node expansion to focus on harder cases. Second, we improve the sample efficiency of BFS through Direct Preference Optimization (DPO) applied to state-tactic pairs automatically annotated with compiler error feedback, refining the LLM’s policy to prioritize productive expansions. Third, we employ length normalization in BFS to encourage exploration of deeper proof paths. BFS-Prover achieves a state-of-the-art score of $72.95\%$ on the MiniF2F test set and therefore challenges the perceived necessity of complex tree search methods, demonstrating that BFS can achieve competitive performance when properly scaled. To facilitate further research and development in this area, we have open-sourced our model at https://huggingface.co/ByteDance-Seed/BFS-Prover-V1-7B.
BFS-Prover: Scalable Best-First Tree Search for LLM-based Automatic Theorem Proving
Ran Xin1,thanks: Key contributors.  Chenguang Xi1,*  Jie Yang2,*  Feng Chen 3,*,thanks: Work done during an internship at ByteDance Seed. Hang Wu2Xia Xiao1Yifan Sun1Shen Zheng1Kai Shen1 1ByteDance Seed 2ByteDance Applied Machine Learning 3Stanford University Correspondence: {ran.xin, shen.kai}@bytedance.com
@@ 1 Introduction
Automatic theorem proving (ATP) in formal languages has recently become a critical benchmark for evaluating the reasoning capabilities of large language models (LLMs). By encoding mathematical problems into formal systems like Lean4, ATP enables the generation of machine-verified proofs for complex mathematical propositions, ensuring logical correctness (Moura and Ullrich, 2021; Polu and Sutskever, 2020). Despite the remarkable success of LLMs in natural language mathematics and reasoning tasks (Lewkowycz et al., 2022; OpenAI, 2023), theorem proving in formal languages presents unique challenges (Yang et al., 2024b; Wu et al., 2022; He et al., 2024; Wang et al., 2023, 2024; Lin et al., 2024; Xin et al., 2024a). Unlike informal reasoning, formal systems require strict adherence to syntax and semantics, as well as the ability to generate valid steps within a highly constrained formal framework. Additionally, the action (tactic) space in ATP is vast, as each proof state can lead to numerous potential tactics, making the search process for valid proofs computationally intensive (Polu et al., 2022).
Tree search algorithms are fundamental to ATP, allowing policy models to navigate large and complex proof space efficiently (Polu and Sutskever, 2020). Among these methods, Monte Carlo Tree Search (MCTS) (Coulom, 2006) has gained popularity due to its ability to balance exploration and exploitation using value functions (critic models) or intrinsic rewards (Browne et al., 2012; Silver et al., 2016). MCTS has demonstrated remarkable success in AlphaZero-style frameworks for games like chess and Go (Silver et al., 2017), where the underlying state spaces are vast but terminal states are well-defined. However, applying MCTS and/or value functions to ATP comes with distinct complications. While games have clear winning and losing conditions, proof search lacks such definitive end states: a proof attempt can theoretically continue indefinitely until a proof or counterexample is found, making it challenging to assess intermediate progress (Han et al., 2021; Lample et al., 2022). Additionally, ATP involves a much larger and more dynamic branching factor, sparse and delayed feedback, and an open-ended reasoning process. These differences highlight the unique demands of ATP and advocate the need for specialized adaptations of search methods to address its complexities.
Best-First Tree Search (BFS) (Pearl, 1984) offers a simpler and more lightweight alternative to MCTS by prioritizing expansions based on accumulated log probabilities from the current node to the root. While its straightforward nature and computational simplicity make it attractive, BFS is often considered suboptimal for theorem proving in the existing literature (Wu et al., 2024a; Li et al., 2024b; Xin et al., 2024b) due to the following hypothesis:
@@ 1.1 Our Contributions
  • Expert iteration with self-filtering. We develop an expert iteration (Anthony et al., 2017) framework that strategically filters out problems solvable by beam search (Steinbiss et al., 1994) node expansions in each round. This filtering is critical as it directs training data accumulation towards harder theorems. As expert iteration progresses, the policy LLM continuously improves, learning a diverse range of tactics and deeper proofs via BFS.
  • Direct preference optimization (DPO) from compiler feedback. We use DPO (Rafailov et al., 2024) to refine the policy LLM by leveraging preference pairs naturally generated during tree search. From a given proof state, each preference pair consists of a positive tactic, which lies on the proof path, and a negative tactic, which leads to a Lean compiler error. DPO sharpens policy’s distribution, enabling it to avoid unproductive tactics and thereby improving the sample efficiency of BFS.
@@
  • Empirical results on MiniF2F. BFS-Prover achieves an accumulative score of 72.95% on the MiniF2F test set, surpassing all state-of-the-art theorem provers in the literature, including DeepSeek-Prover-V1.5 (Xin et al., 2024b), InternLM2.5-StepProver (Wu et al., 2024a), and HunyuanProver (Li et al., 2024b). This result underscores that BFS-Prover can scale to competitive levels in automated theorem proving while maintaining a lightweight design that avoids the complexities of MCTS and value functions.
Roadmap. The remainder of this paper is organized as follows. Section 2 gives an overview of the BFS-Prover system, detailing the expert iteration framework, data filtering, DPO for policy refinement, and length normalization in BFS. Section 3 describes practical implementation and experimental results on the MiniF2F benchmark against leading prover systems. Section 4 concludes the paper.
@@ 2.1 Lean4 Environment and Policy LLM
We adopt LeanDojo (Yang et al., 2024c) as an interactive Python interface for Lean4 integration within BFS-Prover. It transforms Lean4 into a gym-like environment (Brockman, 2016), facilitating interaction between the policy LLM and the formal proof assistant. Specifically, LeanDojo manages state transitions by executing tactics generated by the policy LLM in the Lean4 compiler. If a tactic cannot be executed, LeanDojo captures and returns the corresponding error messages, providing critical feedback for DPO to refine the policy LLM.
@@ 2.3 Expert Iteration
    Direct Preference Optimization (DPO): Instead of re-training the policy LLM using SFT, DPO can be applied as an alternative method to refine the current policy model by leveraging the on-policy Lean error data collected in step 2. Along the proof path, certain generated tactics, expanded from a proof state but not part of the verified proof path, result in Lean compiler errors. These erroneous tactics naturally serve as negative examples when compared to their corresponding valid proof tactics, forming preference tuples. Specifically, for a state $s$ on a proof path, we may form pairs $(a_{w},a_{l})$ where $a_{w}$ is the tactic on the proof path and $a_{l}$ is an error-causing tactic expanded from $s$ if exists; see Fig. 1 for an illustration. The DPO loss $\mathcal{L}_{\text{DPO}}(\theta)$ (Rafailov et al., 2024) is then defined as
    $\mathbb{E}_{(s,a_{w},a_{l})}\!\left[\log\sigma(\beta(r_{\theta}(s,a_{w})-r_{\theta}(s,a_{l})))\right]\!,$
@@ 3 Practical Implementation and Benchmark Results
In this section, we discuss practical implementation details of the BFS-Prover system and present its benchmark results on MiniF2F test. Throughout all experiments, we use Lean 4.7.0.
@@ 3.1 Model, Data, and Training Setup
Base model and initial training data. For illustration purposes, we use Qwen2.5-Math-7B (Yang et al., 2024a) as the base model for fine-tuning the policy LLM in BFS-Prover. In order to properly initialize the expert iteration process, we leverage proof data extracted by LeanDojo (Yang et al., 2024c) from Mathlib (Moura and Ullrich, 2021), which serves as the cold-start dataset. As expert iteration progresses, we further incorporate state-tactic datasets from Lean-Github (Wu et al., 2024b), a compilation of Lean4 repositories from GitHub, and Lean-Workbook (Ying et al., 2024), which focuses on Olympia-level algebra and analysis. These datasets span a diverse range of mathematical topics and formal reasoning tasks, equipping the policy model with the foundational skills necessary for effective tactic generation and proof navigation.
Formal statement corpus. To construct our expert iteration data corpus, we autoformalize the NuminaMath-CoT dataset (Li et al., 2024a) using proprietary tools. We augment this with unproven theorems from Mathlib and formal statements from Lean-Workbook. The resulting corpus comprises approximately 900,000 formal mathematical statements without proofs, providing a comprehensive foundation for expert iteration.
@@
BFS configuration in expert iteration. We set the length normalization parameter $\alpha$ to $0.0$ throughout, to minimize inductive bias in expert iteration. During the beam search filtering stage, we utilize a beam width of $32$ to identify easily solvable theorems. In the subsequent data collection phase, we employ temperature-based sampling with a temperature of $1.0$, nucleus sampling parameter $1.0$, and a sampling width of $2$, 44, or $8$ to explore diverse proof paths.
@@ 3.2 Distributed Best-First Search Infrastructure
  • GPU-Based Policy LLM Pool: Each local machine deploys 8 instances of our 7B policy LLM, each powered by an asynchronous vLLM engine running on a dedicated A100 GPU. These instances form a shared pool that processes concurrent tactic generation requests.
@@ 3.3 Distribution Shift in Expert Iteration
Tactic level. Beyond proof-level evolution, we also observe interesting distributional shifts at the tactic level throughout expert iteration; see Fig. 3. Notably, the policy LLM in the BFS-Prover system maintains a diverse range of tactic lengths without collapsing to a narrow distribution - a common failure mode in reinforcement learning where models tend to converge to a limited set of high-reward actions (Sutton, 2018). Instead, we observe a moderate but meaningful shift from very simple tactics (1-10 tokens) towards more commonly used tactical patterns (11-50 tokens). This shift suggests that through expert iteration, the policy LLM in BFS-Prover learns to generate more sophisticated tactics while maintaining the capability to use simpler ones when appropriate. The preservation of tactic diversity is crucial for effective theorem proving, as different proof states require different levels of tactical complexity, from simple term rewriting to complex algebraic manipulations.
Figure 3: Sample evolution of tactic token length distribution across expert iteration rounds. The plot shows the log-scaled percentage of tactics falling into different token length bins. We have two observations: (1) robustness against mode collapse during expert iteration, evidenced by the maintained diversity across token lengths, and (2) a gradual shift from an emphasis on very short tactics (111010 tokens, decreasing by about $2.3\%$) to a higher prevalence of moderate-length tactics (11115050 tokens, with incremental increases ranging from approximately $1.27\%$ to $0.06\%$). This evolution reflects the improved ability of BFS-Prover to generate more complex, informative tactics while still retaining the capacity to produce concise ones when needed.
@@ 3.4 Results on MiniF2F
In this subsection, we discuss BFS-Prover’s performance on the MiniF2F test benchmark (Zheng et al., 2021) that is a widely recognized dataset for evaluating formal mathematics systems. MiniF2F is consisted of a diverse collection of formalized competition-level mathematics problems. The policy LLM checkpoint for tactic generation used for evaluation was obtained by performing SFT on all state-tactic pairs accumulated through $10$ rounds of expert iteration pipeline of BFS-Prover, followed by an additional round of DPO refinement using Lean error signals as described in Section 2.3.
@@ 3.4.1 Comparison with the State of the Art
We now compare BFS-Prover developed in this work against leading theorem provers in the literature, including DeepSeek-Prover-V1.5 (Xin et al., 2024b), InternLM2.5-StepProver (Wu et al., 2024a), and HunyuanProver (Li et al., 2024b).
@@
Accumulative results. In order to maximize potential proof discovery, we conduct BFS searches under multiple hyperparameter configurations and aggregate the successful proofs. Specifically, we vary the BFS length normalization factor $\alpha$ defined in (1) over $\{0.0,0.5,1.0\}$ to execute different exploration and exploitation trade-offs, while other parameter settings remain unchanged. This diverse search strategy allows BFS-Prover to accumulatively achieve a score of 72.95% on the MiniF2F test, setting a new state-of-the-art result in the formal theorem proving literature. See Fig. 4 for the performance of BFS-Prover when the number of attempts is limited.
@@
DeepSeek-Prover-V1.5 (Xin et al., 2024b) No MCTS $32\times 16\times 400$ 63.5%
@@
InternLM2.5-StepProver (Wu et al., 2024a) Yes BFS $256\times 32\times 600$ 65.9%
HunyuanProver (Li et al., 2024b) Yes BFS $600\times 8\times 400$ 68.4%
@@ 4 Conclusion and Discussion
While the BFS-Prover system demonstrates strong performance in automated theorem proving, several limitations should be acknowledged, particularly regarding model size considerations. Our current implementation relies on a relatively small 7B parameter policy model, which may limit the system’s ability to learn and utilize more sophisticated mathematical reasoning patterns. While larger models (e.g., 32B or 70B parameters) could potentially capture more complex mathematical insights and generate more nuanced tactics, they would introduce significant computational challenges for both training and inference in the context of tree search. This trade-off becomes particularly evident when considering that larger models typically require more GPU memory and have longer inference latency, which could significantly reduce the number of states that can be explored within a fixed time budget. Additionally, complex mathematical proofs can generate extensive state descriptions that may exceed the practical context window of a 7B model, potentially causing the model to miss crucial information needed for generating appropriate tactics.
Appendix: Lean Proofs for IMO Problems Found by BFS-Prover
We present the Lean 4 proofs found by BFS-Prover for several IMO problems in MinF2F test. These examples demonstrate that BFS-Prover can handle complex mathematical reasoning in formal systems, including number theory, inequalities, and geometric relationships. See Fig. 5, 6, 7, 8, and 9 for detailed proofs.
@@
Figure 6: BFS-Prover’s Lean 4 proof of IMO-1964-P2.
@@
    all_goals rw [div_lt_iff (by positivity)] at h2
    all_goals nlinarith [
    sqrt_nonneg (1 + 2 * x), sqrt_nonneg x,
    mul_nonneg h0 (sqrt_nonneg (1 + 2 * x)),
    mul_nonneg h0 (sqrt_nonneg x), mul_self_sqrt h0]